OpenStudy (anonymous):

Help with factoring exponents. :< 5x^2 - 13x + 6. Work with me, don't just give me the answer. I need to know how this works.

sam (.sam.):

Factor the quadratic 5x^2-13x+6. The coefficient of x^2 is 5 and the constant term is 6. The product of 5 and 6 is 30. The factors of 30 which sum to -13 are -3 and -10. So 5x^2-13x+6 = 5x^2-10 x-3x+6 = 5x(x-2)-3(x-2) Then factor x-2 from 5x(x-2)-3(x-2)

OpenStudy (anonymous):

5x - 3?

OpenStudy (radar):

"Factor the quadratic 5x^2-13x+6. The coefficient of x^2 is 5 and the constant term is 6. The product of 5 and 6 is 30. The factors of 30 which sum to -13 are -3 and -10. So 5x^2-13x+6 = 5x^2-10 x-3x+6 = 5x(x-2)-3(x-2) " Looking at the next to last line "So 5x^2 - 13x +6 = 5x^2-10" Why is this so? Assuming this is true, then: